The gaseous hydrocarbon butane, C4H10, burns according to the following equation:
2C4H10(g)+13O2(g)→8CO2(g)+10H2O(g)
How many moles of H2O are produced from the complete reaction of 2.00 mol of C4H10?
Express your answer with the appropriate units.
1
Expert's answer
2015-10-20T04:15:39-0400
2C4H10 (g) + 13O2 (g) → 8CO2 (g) + 10H2O (g)
n(H2O) = 5*n(C4H10) = 5*2 mol = 10 mol

Answer : 10 mol of H2O